OpenStudy (kpop4life123):

What is the length of the missing leg? Round your answer to the nearest tenth.

OpenStudy (anonymous):

use Pythagoras

OpenStudy (kpop4life123):



OpenStudy (anonymous):

a^2 + b^2=c^2

OpenStudy (tanabugg):

31 even though i'll probably get in trouble for telling you answer

OpenStudy (tanabugg):

Was that an answer choice?

OpenStudy (some.random.cool.kid):

yeah might want to watch your back buddy mods might come for you lol

OpenStudy (tanabugg):

30*30= 900 6*6= 36 900 + 36 = 936 square root of 936 = 30.59411... round up to get 31

OpenStudy (tanabugg):

Thanks for the medals!!

OpenStudy (some.random.cool.kid):

now you got the idea! great explanation.

OpenStudy (kpop4life123):

no thank you :)

OpenStudy (tanabugg):

No problem ! This stuff gives me a refresher for pass testing

OpenStudy (kpop4life123):

could u help me on another :)